# Addressing the Devastating Effects of Single-Task Data Poisoning in Exemplar-free Continual Learning

### v1

@@ 3.1 Single Task Data Poisoning: What It Is and Why It Matters
Figure 2: Example of image corruption: Gaussian blur with increasing severity (1-5 from left to right).
@@ 3.2 Data poisoning continual learning with STP
While CIFAR-C dataset is usually used for test-time adaptation scenarios and thus benchmark consists of corrupted test data, we use those predefined transformations to create corrupted images for training the poisoned task. Adapting these predefined transformations results in a clean-label, non-adversarially optimized data poisoning (poisoning based on a pixel distribution change after transformations like e.g. gaussian blur). For examples of corruption severities, see 2. All corruption types are shown in the Appendix 15.
Figure 3: Main parameters of STP attacks: pcp and pn. $pcp$ defines how many classes are poisoned in poisoned task $T_{p}$, $pn$ how many poisons (different corruptions) are used.
3.2.2 Attacks.
@@
  • $pcp$ – poisoned class percentage, defining how many classes are affected by corruptions in $D_{poisoned}$
  • $pn$ – number of poisons (i.e. different corruptions) used to poison data in $D_{poisoned}$
  • $pp$ – the percentage of exemplars poisoned (corrupted) in each class of $D_{poisoned}$
The poisoning attack is changing the distribution of $T_{p}$ data by applying one of the corruptions to $pp$ of images from $pcp$ classes of $D_{clean}$. Let $I_{n}$ be the subset of $n$ indices randomly chosen from $D_{clean}$, $I_{n}\subseteq\{1,2,\ldots,|D_{clean}|\}=I$. Then $D_{p}$ is a set containing $n$ poisoned and $|D_{clean}|-n$ clean samples:
@@
Naming convention for attack evaluation. We evaluate STP with four kinds of attacks:
@@
Figure 5: Difference between clean and poisoned task vectors increase with more layers contributing to task vectors. Consequently, we detect the poison when $\beta>\alpha$.
@@ 4.2 Attacks evaluation
{tabu}
llccc—llllll Row Attack method Dataset p Acc at poisoning time Final Acc
@@
2A) A poisoning singe task in CL sequence decreases performance not only on $T_{p}$, but also on past and future tasks. As expected, the performance decrease is usually the biggest for the poisoned task (see $T_{p}$ columns in Table 4.2). However, Table 4.2 presents further evidence for observation 1): there is almost no difference in accuracy for BEFORE $T_{p}$ and AFTER $T_{p}$ in JOINT training (for rows 2-3 and 5-8 differences between clean and poisoned training in the parenthesis are close to 0), while there is a significant difference when poisoning CL sequence (for rows 10-11, 13-16 differences are negative). At the same time, we see that the relative performance drop is much bigger for tasks before $T_{p}$ than after: overriding patterns to distinguish old classes without access to the data is a bigger problem than struggles in learning new patterns in tasks after $T_{p}$.
@@
Figure 8: Relative accuracy drop for bait ($pn=1$) and multibait ($pn>1$) attacks using different number of poisons. Increasing $pn$ decrease poison effectiveness. Experimental setup: CIFAR100, LWF, Resnet32.
2B) bait attacks affect performance on tasks after $T_{p}$ stronger than base attacks. Figure 8 expands evidence presented in Table 4.2 and show the effects of base and bait attacks with different corruptions on CIFAR100. Interestingly, bait attacks are stronger than base while poisoning much less exemplars of $T_{p}$. base attacks primarily reduce accuracy on past tasks, affecting model stability, while bait attacks decrease performance on both past and future tasks. We hypothesize that spurious correlations (’baits’) in $T_{p}$ data impair future task learning more than single corruption-based data distribution shifts in base attacks.
3) multibase and multibait attacks are less effective than their counterparts with single corruption. Comparing results for multi-corruption attacks and single corruption attacks in Table 4.2 we see a bigger performance decrease for the latter ones: impact of one corruption is stronger than multiple, when we poison constant number of classes (with constant $pcp$). Results presented on Figure 8 support this observation. The bigger the number of used corruptions, the smaller the poisoning effect on both past and future tasks.
@@ A.1 Extended investigation of STP Attacks
8) Poisoning a vast percentage of class exemplars is more important than corruption severity. Figure 14 is an extension of Figure 9 presented in the main body of this work. It shows (in accordance to observations stated in 4)) that two conditions are necessary for a successful STP attack. Firstly, the corruption of the image itself cannot be negligible. Very low severity does not disrupt the training (4A). Secondly, the vast percentage of $T_{p}$ must be poisoned (4B). From those two conditions, the latter has greater significance. When poisoning $>=90\%$ of class exemplars in $T_{p}$ performance decrease even with smaller severities of corruption. On the other hand, when less exemplars are poisoned the accuracy drop is negligible even for the strongest severity considered.
@@
9A) STP attacks are successful for streams with more complex images. To confirm that STP attacks pose a threat also for more complex CL benchmarks, we conduct experiments on Split Tiny Imagenet. Table A.1 shows similar results to those presented in main part of this work (see Table 4.2): in terms of FINAL ACC, the most affected is accuracy on poisoned task ($T_{p}$), but there exists also significant drop for $BEFORE~{}T_{p}$ and $AFTER~{}T_{p}$ results. The results further reinforce the evidence supporting the effectiveness of STP attacks and validate their impact on more complex datasets beyond CIFAR-10/100.
9B) base attacks are stronger with increasing number of classes in $T_{p}$ and increasing classification complexity. Table A.1 shows strong performance drop for base  attack. We attribute it to two following factors: 1) the number of classes in $T_{p}$, 2) total number of classes (i.e. classification task complexity). Regarding the TOTAL FINAL ACC, the decreases become more pronounced as the number of classes and dataset complexity increase: -2.7 (with $T_{p}$ having two classes for CIFAR-10), -9.7 (with $T_{p}$ having 10 classes for CIFAR-100), and -11.6 (with $T_{p}$ having 20 classes for Tiny ImageNet).
10) Performance on past and future tasks is not affected for STP attacks on memory-based CL methods. In the Limitations section of this work, we note that the STP framework is effective for investigating memory-free CL methods but less applicable to those utilizing exemplars. The STP framework restricts the adversary to poisoning only a single task. When only one task is poisoned, CL methods with memory maintain an unrealistic guarantee that all previous samples stored in the buffer are clean during the $T_{p}$ task (since $T_{p}$ is the only poisoned task). Introducing an additional buffer of clean samples undermines the concept of ’learning the task in isolation’ and mitigates the STP attack, thereby creating a scenario more akin to joint training. Results for simple finetuning with buffer and GDUMB (Prabhu et al. (2020)) presented in Table A.1 confirm that (similarly to JOINT training) attack decrease only performance on $T_{p}$, while there are no accuracy drops for $BEFORE~{}T_{p}$ and $AFTER~{}T_{p}$ results (numbers in parenthesis are positive or close to zero).
Additional results for LwM. Table A.1 presents STP attacks evaluation for additional data regualrization method – LwM (Dhar et al. (2019)). These results are consistent with observations described in the main part of this work.
@@ Appendix B Experimental Setup details
Corruptions Figure 15 shows a set of CIFAR-C based corruptions used to poison data in this work. While CIFAR-C dataset is usually used for test-time adaptation scenarios and thus benchmark consists of corrupted test data, we use those predefined transformations to create corrupted images for training the poisoned task. For all results with single poison reported in tables we use $gaussian\_blur$ corruption (see Figure 2). For attacks with higher number of poisons (multibase and multibait) we use five poisons: $gaussian\_blur$, $gaussian\_noise$, $contrast$, $pixelate$, $jpeg\_compression$. We show attacks using all kinds of corruptions (each as a single poison for base or bait attacks) on Figure 8.
@@
Additional parameters. We use lambda=10 (distillation loss parameter) for both LWM and LwF methods. For additional experiments on LwM (see Table A.1) we use additionally gamma=20 (attention loss parameter). We use data augmentations from FeTrIL (Petit et al. (2023)).

### v2

@@ 3.1 Single Task Data Poisoning: What It Is and Why It Matters
Figure 2: Image corruption example: Gaussian blur with increasing severity (1–5).
@@ 3.2 Data poisoning continual learning with STP
While CIFAR-C dataset is usually used for test-time adaptation scenarios and thus benchmark consists of corrupted test data, we use those predefined transformations to create corrupted images for training the poisoned task. Adapting these predefined transformations results in a clean-label, non-adversarially optimized data poisoning (poisoning based on a pixel distribution change after transformations like e.g. gaussian blur). For examples of corruption severities, see Figure 2. All corruption types are shown in the Appendix 15.
@@
  • pcppcp – poisoned class percentage, defining how many classes are affected by corruptions in $D_{poisoned}$
  • pnpn – number of poisons (i.e. different corruptions) used to poison data in $D_{poisoned}$
  • pppp – the percentage of exemplars poisoned (corrupted) in each class of $D_{poisoned}$
The poisoning attack is changing the distribution of $T_{p}$ data by applying one of the corruptions to $pp$ % of images from pcppcp classes of $D_{clean}$. Let $I_{n}$ be the subset of $n$ indices randomly chosen from $D_{clean}$, $I_{n}\subseteq\{1,2,\ldots,|D_{clean}|\}=I$. Then $D_{p}$ is a set containing $n$ poisoned and $|D_{clean}|-n$ clean samples:
@@
3.2.2 Attacks.
Figure 3: Main parameters of STP attacks: pcppcp defines how many classes are poisoned in the poisoned task $T_{p}$, pnpn how many poisons (corruptions) are used.
Naming convention for attack evaluation.
We evaluate STP with four kinds of attacks:
@@
Figure 5: Difference between clean and poisoned task vectors increases with more layers contributing to task vectors. Consequently, we detect the poison when $\beta>\alpha$.
@@ 4.2 Attacks evaluation
Row Attack method Dataset p Acc at poisoning time Final Acc
@@
2A) A poisoning singe task in CL sequence decreases performance not only on $T_{p}$, but also on past and future tasks. As expected, the performance decrease is usually the biggest for the poisoned task (see $T_{p}$ columns in Table 2). However, Table 2 presents further evidence for observation 1): there is almost no difference in accuracy for BEFORE $T_{p}$ and AFTER $T_{p}$ in JOINT training (for rows 2-3 and 5-8 differences between clean and poisoned training in the parenthesis are close to 0), while there is a significant difference when poisoning CL sequence (for rows 10-11, 13-16 differences are negative). At the same time, we see that the relative performance drop is much bigger for tasks before $T_{p}$ than after: overriding patterns to distinguish old classes without access to the data is a bigger problem than struggles in learning new patterns in tasks after $T_{p}$.
@@
Figure 8: Relative accuracy drop for bait ($pn=1$) and multibait ($pn>1$) attacks using different number of poisons. Increasing pnpn decrease poison effectiveness. Experimental setup: CIFAR100, LWF, Resnet32.
2B) bait attacks affect performance on tasks after $T_{p}$ stronger than base attacks. Figure 8 expands evidence presented in Table 2 and show the effects of base and bait attacks with different corruptions on CIFAR100. Interestingly, bait attacks are stronger than base while poisoning much less exemplars of $T_{p}$. base attacks primarily reduce accuracy on past tasks, affecting model stability, while bait attacks decrease performance on both past and future tasks. We hypothesize that spurious correlations (’baits’) in $T_{p}$ data impair future task learning more than single corruption-based data distribution shifts in base attacks.
3) multibase and multibait attacks are less effective than their counterparts with single corruption. Comparing results for multi-corruption attacks and single corruption attacks in Table 2 we see a bigger performance decrease for the latter ones: impact of one corruption is stronger than multiple, when we poison constant number of classes (with constant pcppcp). Results presented on Figure 8 support this observation. The bigger the number of used corruptions, the smaller the poisoning effect on both past and future tasks.
@@ A.1 Extended investigation of STP Attacks
8) Poisoning a vast percentage of class exemplars is more important than corruption severity. Figure 14 is an extension of Figure 9 presented in the main body of this work. It shows (in accordance to observations stated in 4)) that two conditions are necessary for a successful STP attack. Firstly, the corruption of the image itself cannot be negligible. Very low severity does not disrupt the training (4A). Secondly, the vast percentage of $T_{p}$ must be poisoned (4B). From those two conditions, the latter has greater significance. When poisoning $90\%$ of class exemplars in $T_{p}$ performance decrease even with smaller severities of corruption. On the other hand, when less exemplars are poisoned the accuracy drop is negligible even for the strongest severity considered.
@@
9A) STP attacks are successful for streams with more complex images. To confirm that STP attacks pose a threat also for more complex CL benchmarks, we conduct experiments on Split Tiny Imagenet. Table 3 shows similar results to those presented in main part of this work (see Table 2): in terms of FINAL ACC, the most affected is accuracy on poisoned task ($T_{p}$), but there exists also significant drop for $BEFORE~T_{p}$ and $AFTER~T_{p}$ results. The results further reinforce the evidence supporting the effectiveness of STP attacks and validate their impact on more complex datasets beyond CIFAR-10/100.
9B) base attacks are stronger with increasing number of classes in $T_{p}$ and increasing classification complexity. Table 3 shows strong performance drop for base  attack. We attribute it to two following factors: 1) the number of classes in $T_{p}$, 2) total number of classes (i.e. classification task complexity). Regarding the TOTAL FINAL ACC, the decreases become more pronounced as the number of classes and dataset complexity increase: -2.7 (with $T_{p}$ having two classes for CIFAR-10), -9.7 (with $T_{p}$ having 10 classes for CIFAR-100), and -11.6 (with $T_{p}$ having 20 classes for Tiny ImageNet).
10) Performance on past and future tasks is not affected for STP attacks on memory-based CL methods. In the Limitations section of this work, we note that the STP framework is effective for investigating memory-free CL methods but less applicable to those utilizing exemplars. The STP framework restricts the adversary to poisoning only a single task. When only one task is poisoned, CL methods with memory maintain an unrealistic guarantee that all previous samples stored in the buffer are clean during the $T_{p}$ task (since $T_{p}$ is the only poisoned task). Introducing an additional buffer of clean samples undermines the concept of ’learning the task in isolation’ and mitigates the STP attack, thereby creating a scenario more akin to joint training. Results for simple finetuning with buffer and GDUMB (Prabhu et al. (2020)) presented in Table 6 confirm that (similarly to JOINT training) attack decrease only performance on $T_{p}$, while there are no accuracy drops for $BEFORE~T_{p}$ and $AFTER~T_{p}$ results (numbers in parenthesis are positive or close to zero).
Additional results for LwM. Table 7 presents STP attacks evaluation for additional data regualrization method – LwM (Dhar et al. (2019)). These results are consistent with observations described in the main part of this work.
@@
A.3 Additional analysis and t-SNE visualizations of STP.
(a) clean
(b) bait attack
Figure 16: Confusion matrices for CIFAR10 classes at the end of the CL training. Clean run (left) consequently has high values on the diagonal, with low number of prediction errors. bait attack (right) on the 2nd task results in almost no predictions for class with corrupted images (class 4, in red), while class 5 has very high number of wrong predictions.
bait affects classes in the poisoned task unequally. As shown in Figure 16, confusion matrices at the end of training reveal class-specific effects of the bait attack. In Task 3, where two classes are involved, the class with label 4 (corrupted images) is almost never predicted – its diagonal entry is effectively zero. Conversely, class 5 (uncorrupted) is over-predicted, reducing overall accuracy for other classes. This indicates that the attack’s effect depends strongly on which class receives the corruption.
(a) After 2nd task
(b) After 3rd task
(c) After 4th task
(d) After 2nd task
(e) After 3rd task
(f) After 4th task
Figure 17: t-SNE visualization of the model output space on clean test data. 1) when second task is clean (upper row, (a)-(c)) 2) when second task is poisoned using bait attack (bottom row, (d)-(f)).
Model representation does not collapse after poisoning: t-SNE visualizations. We reject the hypothesis that the model collapses to a single representation following the poisoned task. Instead, the baitattack introduces a subtle distribution shift that causes misclassification of clean examples. This is visible when comparing (a) and (d) in Figure17, or the first row in Figure 18. These initial misalignments propagate into future tasks, slightly distorting class boundaries compared to the clean setup.
Figure 18: t-SNE visualization of the model output space on clean test data: 1) when all tasks are clean (left) 2) when the first task is poisoned using bait attack (right).
@@ Appendix B Experimental Setup details
Corruptions Figure 15 shows a set of CIFAR-C based corruptions used to poison data in this work. While CIFAR-C dataset is usually used for test-time adaptation scenarios and thus benchmark consists of corrupted test data, we use those predefined transformations to create corrupted images for training the poisoned task. For all results with single poison reported in tables we use gaussian_blurgaussian\_blur corruption (see Figure 2). For attacks with higher number of poisons (multibase and multibait) we use five poisons: gaussian_blurgaussian\_blur, gaussian_noisegaussian\_noise, contrastcontrast, pixelatepixelate, jpeg_compressionjpeg\_compression. We show attacks using all kinds of corruptions (each as a single poison for base or bait attacks) on Figure 8.
@@
Additional parameters. We use lambda=10 (distillation loss parameter) for both LWM and LwF methods. For additional experiments on LwM (see Table 7) we use additionally gamma=20 (attention loss parameter). We use data augmentations from FeTrIL (Petit et al. (2023)).